# Steiner Traveling Salesman Problem with Time Windows and Pickup–Delivery: integrating classical and quantum optimization

### v2

@@ 4.3.1 Equivalence Between ABF and NBF Formulations
The numerical results comparing the two formulations are presented in subsection 6.4.1, where we assess their performance and verify the correspondence of optimal solutions.
@@ 6.1 Computational setup and solver configuration
The CQM formulation is implemented using D-Wave’s Ocean SDK, specifically dimod version 0.12.18 and dwave-system version 1.28.0. All problem instances are solved through D-Wave’s hybrid solver service for CQMs, identified as hybrid_constrained_quadratic_model_version1p. We impose a time_limit of 5 seconds.
In practice, the CQM framework automatically converts each constraint into an energy-contributing term within the Hamiltonian by means of an internal penalty calibration mechanism. This automated process determines appropriate penalty scales that balance constraint satisfaction and objective minimization, ensuring that feasible solutions are prioritized without introducing numerical instability. As a result, users can focus on expressing the optimization model (defining variables, constraints, and objectives) without manually managing penalty coefficients. In our implementation, all decision variables are represented as binary or integer types, and both the objective and constraints are formulated using D-Wave’s dimod.ConstrainedQuadraticModel() interface, as documented in the official D-Wave Ocean SDK manual [DWaveCQMdocs]. The completed model is automatically compiled and submitted to the hybrid solver, which orchestrates the interplay between classical and quantum resources to enforce constraints and minimize the Hamiltonian energy.
@@
We utilize the LeapCQMHybrid solver, a proprietary hybrid optimization tool developed by D-Wave Systemsa cloud-based framework designed to integrate classical and quantum computing resources within a unified optimization pipeline. As described in recent works on hybrid architectures [osaba2025d, bertuzzi2024evaluation, willsch2022benchmarking], the Hybrid Solver Service (HSS) follows a modular asynchronous workflow that combines the strengths of classical heuristics with quantum annealing-based exploration. From a structural perspective, each instance submitted to the hybrid solver is decomposed into multiple branches, which are executed in parallel on independent computational threads. Each branch contains two tightly coupled modules:
@@
Figure 1: General structure of D-Wave’s hybrid solvers, adapted from [osaba2025d]. Each branch includes a CM and a QM that interact asynchronously to refine candidate solutions.
The proportion of computation devoted to classical versus quantum phases is dynamically determined by the solver’s internal scheduler and is not user-controllable. According to D-Wave’s technical documentation [DWave2022performance] the frequency of quantum queries depends on problem size, constraint density, and classical progress metrics. In typical workloads, quantum queries account for a minority of total iterations but play a crucial exploratory role, guiding the search toward low-energy basins that are difficult to reach through purely classical heuristics. Although the internal configuration of the hybrid solver is proprietary, its design philosophy follows that of D-Wave’s hybrid architecture, where quantum and classical resources cooperate asynchronously. The classical component ensures scalability and constraint management, while the quantum component contributes through stochastic sampling based on quantum tunneling. This mechanism allows the solver to explore low-energy regions of the search space that are typically difficult to reach with classical heuristics alone. The use of the quantum module is therefore not aimed at achieving quantum speedup in a strict complexity-theoretic sense, but at improving diversification and refinement within the hybrid optimization loop. As evidenced in recent benchmark studies [osaba2025d, bertuzzi2024evaluation, willsch2022benchmarking], such hybridization can enhance convergence reliability and solution quality for large combinatorial problems that exceed the capacity of current QPUs.
@@
In our experiments, quantum queries are executed on D-Wave’s Advantage2_system1.4 QPU, the most advanced QPU architecture currently available. This system consists of 4596 qubits arranged according to a Zephyr connectivity graph. Solver parameters are kept at their default values, with the only tunable parameter, time_limit, left unchanged across all instances to ensure consistency and reproducibility. Due to the proprietary nature of LeapCQMHybrid, detailed information about its internal quantum routines, resource allocation, and the exact number of qubits used per instance is not disclosed. For further technical insights and practical applications, readers are referred to D-Wave’s official documentation and recent literature on hybrid quantum computing [osaba2025d, bertuzzi2024evaluation, willsch2022benchmarking].
@@ 6.2 Generation of instances
The node coordinates are generated by adapting the procedure proposed by Letchford et al. (2013)[letchford2013compact], a widely adopted benchmark method for constructing instances of the STSP. Specifically, we consider a set of $n$ vertices $V=\{v_{1},v_{2},\dots,v_{n}\}$ uniformly positioned along the circumference of a circle with fixed radius. Each vertex $v_{i}$ is assigned a Cartesian coordinate $(x_{i},y_{i})$ computed as:
@@
where $R$ is the radius of the circle ($R=100$) and $i\in\{0,\dots,n-1\}$. From this layout, we consider all possible undirected edges between vertices (excluding the depot) and compute the Euclidean distance between endpoints as the base length of each arc. The candidate arcs are then sorted in ascending order of their length. Edges are incrementally added to the graph while ensuring the following constraints are satisfied:
@@
    The new edge does not intersect any of the previously selected edges;
@@
    The angle between any two adjacent edges incident to the same vertex is not less than $60^{\circ}$.
This selective insertion of edges yields a sparse, planar-like structure while preserving geometric properties inspired by real-world networks. To ensure connectivity, additional arcs are inserted where necessary. Specifically, for any vertex without incoming or outgoing arcs, at least one arc is artificially added to/from a randomly chosen adjacent vertex, subject to the same spatial consistency constraints. These additional arcs are assigned both a length equal to their Euclidean distance. Finally, all duplicate arcs, i.e., arcs defined on the same pair of nodes $(i,j)$, are removed to avoid redundancy and inconsistencies in the graph structure. In accordance with the instance generation procedure described by Letchford et al. [letchford2013compact], each node is independently designated as a required node with probability $p=0.7$, and one of the required nodes is randomly selected as the depot. This intermediate configuration, corresponding to approximately two-thirds of the nodes, yields instances of balanced density and computational complexity, in line with standard STSP benchmark practices.
For the parameters related to time windows, service times, demand, and vehicle capacity, we first carried out a literature review and identified Li and Lim (2001)[li2001metaheuristic] as a standard reference in the study of VRP-TWPDP.
We develop a custom instance generator specifically designed for the STSP-TWPD. The parameter generation follows a randomized procedure with a fixed seed to ensure reproducibility. The updated generation process is as follows:
@@
    Time Windows ($a_{i}$, $b_{i}$): for each required node ($i\in V_{r}$), time windows are constructed based on both the shortest and longest path distances from the depot:
    • (a)
      The shortest paths $l_{\min}[i]$ are computed using Dijkstra’s algorithm and the longest paths $l_{\max}[i]$ via a Bellman-Ford’s strategy adapted for our graphs.
    • (b)
      Two weighted target values are used to introduce variability in the generation of time windows, ensuring a more diverse and realistic distribution:
      $l^{(1)}_{i}$ $=\alpha\cdot l_{\min}[i]+(1-\alpha)\cdot l_{\max}[i],$
      $l^{(2)}_{i}$ $=\beta\cdot l_{\min}[i]+(1-\beta)\cdot l_{\max}[i],$
      where $\alpha=0.7$ and $\beta=0.3$.
    • (c)
      The set of nodes $V_{r}$ is divided into two equal groups for generating early time windows:
      • i.
        For 50% of the nodes in $V_{r}$, the early window start time $a_{i}$ is generated as
@@
        where $\epsilon_{i}$ is a random value uniformly drawn from the interval $(0,\,0.5\cdot l^{(1)}_{i})$.
      • ii.
        For the remaining 50%, early windows are generated as
@@
        with $\epsilon_{i}$ is a random value taken from the interval $(0,\,0.5\cdot l^{(2)}_{i})$.
    • (d)
      The latest service start time $b_{i}$ is determined by adding a random value in $(l^{(2)}_{i},\;l^{(1)}_{i})$ to $a_{i}$, ensuring that $b_{i}\geq a_{i}$:
      $b_{i}=a_{i}+w_{i}$
    For non-required nodes ($i\in V\setminus V_{r}$), we fix $a_{i}=0$ and $b_{i}=10^{5}$.
  • 2.
    Service time ($s_{i}$): each required node receives a duration $s_{i}$ random in $\{5,\;10\}$. For all $i\in V\setminus V_{r}$, $s_{i}=0$.
  • 3.
    Demands ($d_{i}$): the nodes in the set $Vr$ are divided into two equal parts:
    • (a)
      50% of the nodes are assigned a positive demand $d_{i}\in[10,20]$, representing pickups.
    • (b)
      The remaining 50% are assigned a negative demand $d_{i}\in[-5,-1]$, representing deliveries.
    All nodes in $V\setminus V_{r}$ have zero demand.
  • 4.
    Vehicle Capacity ($Q$): the capacity is set to match the total absolute demand, with a small random offset:
@@
  • 5.
    Big-M Constant ($M$): the constant $M$ is defined to upper-bound any feasible arrival time:
@@
    By summing these values, $M$ provides a safe upper bound on the arrival time at any node in any feasible route.
Table 4 summarizes the generation rules and notation.
@@
$V_{r}$ Subset of $V$ consisting of 70% of the nodes (required nodes)
AA Set of arcs derived from a sparse directed graph
$l_{\min}[i]$ Shortest path from depot to $i$ (Dijkstra’s algorithm)
$l_{\max}[i]$ Longest path from depot to $i$ (Bellman-Ford’s algorithm)
$a_{i}$ $\forall i\in V_{r}$: $a_{i}=t^{(1)}_{i}+\epsilon_{i}$ or $t^{(2)}_{i}+\epsilon_{i}$
@@
$b_{i}$ $\forall i\in V_{r}$: $b_{i}=a_{i}+w_{i}$, with $w_{i}$ random in $(l^{(2)}_{i},l^{(1)}_{i})$
$\forall i\in V\setminus V_{r}$: $b_{i}=100000$
$s_{i}$ $\forall i\in V_{r}$: $s_{i}$ random in $\{5,\dots,10\}$
@@
$d_{i}$ 50% of nodes in $V_{r}$: $d_{i}$ random in $\{10,\dots,20\}$ (pickups)
50% of nodes in $V_{r}$: $d_{i}$ random in $\{-5,\dots,-1\}$ (deliveries)
@@
$l_{ij}$ Euclidean distance between each pair of nodes $(i,j)$
$l_{k}$ Travel time of arc $k\in A$
$Q$ $Q=\sum_{i\in V}|d_{i}|+\theta$, $\theta$ random in $(\min_{i}|d_{i}|,\;\max_{i}d_{i})$
$M$ $M=\max_{i\in V}b_{i}+\max_{i\in V}s_{i}+\max_{(i,j)\in A}l_{ij}$
Table 4: Parameters and generation method for random STSP-TWPD instances.

### v3

@@ 4.3.1 Equivalence Between ABF and NBF Formulations
The numerical results comparing the two formulations are presented in subsection 6.4, where we assess their performance and verify the correspondence of optimal solutions.
@@ 6.1 Computational setup and solver configuration
The CQM formulation is implemented using D-Wave’s Ocean SDK, specifically dimod version 0.12.18 and dwave-system version 1.28.0. All problem instances are solved through D-Wave’s hybrid solver service for CQMs, identified as hybrid_constrained_quadratic_model_version1p. We impose a time_limit of 5 seconds. The CQM framework represents a generalization of the traditional QUBO approach, designed to support optimization problems that include both objective functions and explicit constraints. Unlike QUBO models, where all constraints must be manually embedded into the objective function through penalty terms with user-defined coefficients, the CQM formulation allows equality and inequality constraints to be declared directly within the model. This provides a higher level of abstraction and avoids the empirical and problem-dependent process of penalty weight tuning, which often affects convergence and solution quality.
In practice, the CQM framework automatically converts each constraint into an energy-contributing term within the Hamiltonian by means of an internal penalty calibration mechanism. This automated process determines appropriate penalty scales that balance constraint satisfaction and objective minimization, ensuring that feasible solutions are prioritized without introducing numerical instability. As a result, users can focus on expressing the optimization model (defining variables, constraints, and objectives) without manually managing penalty coefficients. In our implementation, all decision variables are represented as binary or integer types, and both the objective and constraints are formulated using D-Wave’s dimod.ConstrainedQuadraticModel() interface, as documented in the official D-Wave Ocean SDK manual [13]. The completed model is automatically compiled and submitted to the hybrid solver, which orchestrates the interplay between classical and quantum resources to enforce constraints and minimize the Hamiltonian energy.
@@
We utilize the LeapCQMHybrid solver, a proprietary hybrid optimization tool developed by D-Wave Systems,to solve the CQM models defined for the problems considered in this work. This solver is a cloud-based framework designed to integrate classical and quantum computing resources within a unified optimization pipeline. As described in recent works on hybrid architectures [36, 3, 49], the Hybrid Solver Service (HSS) follows a modular asynchronous workflow that combines the strengths of classical heuristics with quantum annealing-based exploration. From a structural perspective, each instance submitted to the hybrid solver is decomposed into multiple branches, which are executed in parallel on independent computational threads. Each branch contains two tightly coupled modules:
@@
Figure 1: General structure of D-Wave’s hybrid solvers, adapted from [36]. Each branch includes a CM and a QM that interact asynchronously to refine candidate solutions.
The proportion of computation devoted to classical versus quantum phases is dynamically determined by the solver’s internal scheduler and is not user-controllable. According to D-Wave’s technical documentation [33] the frequency of quantum queries depends on problem size, constraint density, and classical progress metrics. In typical workloads, quantum queries account for a minority of total iterations but play a crucial exploratory role, guiding the search toward low-energy basins that are difficult to reach through purely classical heuristics. Although the internal configuration of the hybrid solver is proprietary, its design philosophy follows that of D-Wave’s hybrid architecture, where quantum and classical resources cooperate asynchronously. The classical component ensures scalability and constraint management, while the quantum component contributes through stochastic sampling based on quantum tunneling. This mechanism allows the solver to explore low-energy regions of the search space that are typically difficult to reach with classical heuristics alone. The use of the quantum module is therefore not aimed at achieving quantum speedup in a strict complexity-theoretic sense, but at improving diversification and refinement within the hybrid optimization loop. As evidenced in recent benchmark studies [36, 3, 49], such hybridization can enhance convergence reliability and solution quality for large combinatorial problems that exceed the capacity of current QPUs.
@@
In our experiments, quantum queries are executed on D-Wave’s Advantage2_system1.4 QPU, the most advanced QPU architecture currently available. This system consists of 4596 qubits arranged according to a Zephyr connectivity graph. Solver parameters are kept at their default values, with the only tunable parameter, time_limit, left unchanged across all instances to ensure consistency and reproducibility. Due to the proprietary nature of LeapCQMHybrid, detailed information about its internal quantum routines, resource allocation, and the exact number of qubits used per instance is not disclosed. For further technical insights and practical applications, readers are referred to D-Wave’s official documentation and recent literature on hybrid quantum computing [36, 3, 49].
@@ 6.2 Generation of instances
The generation of test instances plays a critical role in assessing the robustness and generalizability of the proposed formulations. This subsection describes the procedure adopted to construct the benchmark instances used in the computational study. In particular, we detail how the graph topology is generated and how all problem components, including time windows, service times, pickup–delivery demands, and vehicle capacity, are defined. The overall process is designed to produce different, controlled, and fully reproducible instances. For clarity, we summarize the instance generation procedure in a step-by-step manner:
  1. 1.
    Construction of the underlying graph structure;
  2. 2.
    Selection of required nodes;
  3. 3.
    Generation of travel times based on Euclidean distances;
  4. 4.
    Generation of time windows;
  5. 5.
    Assignment of service times;
  6. 6.
    Generation of pickup and delivery demands;
  7. 7.
    Definition of vehicle capacity;
  8. 8.
    Definition of auxiliary parameters (e.g., Big-M constant).
The design of the instance generator is guided by three main principles. First, realism: the generated instances aim to reflect key characteristics of real-world routing problems, including heterogeneous time windows, service times, and mixed pickup–delivery operations. Second, controllability: the generation process allows us to systematically vary structural and operational parameters (e.g., graph sparsity and proportion of required nodes) in order to analyze their impact on problem difficulty. Third, reproducibility: all random components are generated using a fixed seed, ensuring that the instances can be exactly replicated.
Graph construction
The node coordinates are generated by adapting the procedure proposed by Letchford et al. [29], a widely adopted benchmark method for constructing instances of the STSP. Specifically, we consider a set of $n$ vertices $V=\{v_{1},v_{2},\dots,v_{n}\}$ uniformly positioned along the circumference of a circle with fixed radius. Each vertex $v_{i}$ is assigned a Cartesian coordinate $(x_{i},y_{i})$ computed as:
@@
where $R=100$ and $i\in\{0,\dots,n-1\}$. From this layout, we consider all possible undirected edges between vertices and compute the Euclidean distance between endpoints as the base length of each arc. The candidate edges are sorted in ascending order of their length and incrementally added to the graph. Following Letchford et al. [29], edges are inserted while enforcing geometric constraints to preserve a sparse and planar-like structure. In particular, each edge is accepted only if:
@@
    it does not intersect any previously selected edge;
@@
    the angle between any two adjacent edges incident to the same vertex is not smaller than a given threshold $\theta$.
In the benchmark proposed by [29], an angle of $\theta=60^{\circ}$ is imposed between any two adjacent edges incident to the same vertex. We extend the generation scheme by considering an additional angular threshold $\theta=80^{\circ}$. This allows us to generate graph topologies with different levels of sparsity. In particular, larger values of $\theta$ restrict the number of feasible edges, resulting in sparser and more constrained network structures. To ensure connectivity, additional arcs are inserted when necessary. Specifically, for any vertex without incident arcs, at least one arc is added by connecting it to a nearby vertex, while preserving the geometric consistency of the graph. These arcs are assigned a cost equal to their Euclidean distance. Duplicate arcs are removed to avoid redundancy.
The set of required nodes $V_{r}\subseteq V$ is determined by selecting a fixed proportion of nodes. While the [29] benchmark considers a single ratio $|V_{r}|/|V|=0.3$, we broaden this setting by considering $|V_{r}|/|V|\in{0.2,0.3,0.4,0.5,0.6}$, thereby enabling a more comprehensive analysis of how the proportion of required nodes affects the problem structure and solution characteristics.
The joint variation of the angular constraint and the required-node ratio defines a structured family of instance classes, summarized in Table 4. These parameters jointly control the structural properties of the graph. In particular, larger angular thresholds produce sparser and more constrained networks, while higher values of $|V_{r}|/|V|$ increase the number of mandatory visits, leading to denser routing requirements.
The resulting classes are grouped into two families: A1–A4 for $\theta=60^{\circ}$ and B1–B5 for $\theta=80^{\circ}$. Compared to the baseline instances, this extended design enables a systematic investigation of the interplay between graph sparsity and the density of required nodes.
Class Angle constraint $|V_{r}|/|V|$ Topology effect
A1 $60^{\circ}$ 0.20 Very sparse
A2 ([29]’s instances) $60^{\circ}$ 0.30 Moderate sparsity
A3 $60^{\circ}$ 0.40 Sparse
A4 $60^{\circ}$ 0.50 Moderate density
A5 $60^{\circ}$ 0.60 Dense
B1 $80^{\circ}$ 0.20 Very sparse
B2 $80^{\circ}$ 0.30 Sparse
B3 $80^{\circ}$ 0.40 Sparse
B4 $80^{\circ}$ 0.50 Highly constrained
B5 $80^{\circ}$ 0.60 Very highly constrained
Table 4: Instance classes obtained by varying the angular constraint and the proportion of required nodes.
To the best of our knowledge, no benchmark generation procedure currently exists for the STSP-TWPD, nor for its individual variants STSP-TW and STSP-PD. To fill this gap, we propose a novel instance generation framework specifically tailored to these problem settings. The proposed generator combines graph-based structural information with controlled randomization to produce realistic and diverse instances, while ensuring full reproducibility through the use of a fixed random seed. This design allows us to systematically control the temporal and operational characteristics of the problem, enabling a meaningful and comprehensive computational evaluation.
Time windows
Let $l_{\min}[i]$ denote the shortest path distance from the depot to node $i$, computed via Dijkstra’s algorithm, and let $l_{\max}[i]$ denote a longest-path estimate obtained through a Bellman-Ford-based procedure. For each node $i\in V_{r}$, two reference values $l^{(1)}_{i}$ and $l^{(2)}_{i}$ are derived by combining $l_{\min}[i]$ and $l_{\max}[i]$ as to produce two distinct reference levels, with $l^{(1)}_{i}\leq l^{(2)}_{i}$. The set of required nodes $V_{r}$ is partitioned into two equally sized subsets for the generation of early time windows.
@@
    For 50% of the nodes in $V_{r}$, the earliest service start time $a_{i}$ is defined as
@@
    where $\epsilon_{i}$ is a random value uniformly distributed in the interval $(0,\;0.5\cdot l^{(1)}_{i})$.
  • 2.
    For the remaining 50% of nodes, the earliest service start time is generated as
@@
    where $\epsilon_{i}$ is uniformly drawn from $(0,\;0.5\cdot l^{(2)}_{i})$.
The latest service start time $b_{i}$ is then computed by adding a random value $w_{i}$ to $a_{i}$, where $w_{i}$ is a random offset, uniformly drawn from the interval $(l^{(1)}_{i},\;l^{(2)}_{i})$, ensuring that $b_{i}\geq a_{i}$:
$b_{i}=a_{i}+w_{i}.$
For non-required nodes $i\in V\setminus V_{r}$, we set $a_{i}=0$ and $b_{i}=10^{5}$.
Service times
For each $i\in V_{r}$, the service time is sampled as $s_{i}\in\{5,10\}$. For all $i\in V\setminus V_{r}$, $s_{i}=0$.
Demands
The set of required nodes is partitioned into two equally sized subsets, corresponding to pickup and delivery operations. For the first subset (pickup nodes), positive demands are assigned by sampling uniformly from the interval $\{10,\dots,20\}$. For the second subset (delivery nodes), negative demands are generated uniformly from $\{-5,\dots,-1\}$. All non-required nodes are assigned zero demand.
Vehicle capacity
The capacity is set to match the total absolute demand, with a small random offset:
@@
Big-M constant
The constant $M$ is defined to upper-bound any feasible arrival time:
@@
By summing these values, $M$ provides a safe upper bound on the arrival time at any node in any feasible route. Table 5 summarizes the generation rules and notation.
Table 5: Parameters and generation rules for STSP-TWPD instances.
@@
$V_{r}$ Subset of required nodes, with $|V_{r}|/|V|\in\{0.2,0.3,0.4,0.5,0.6\}$
AA Set of arcs derived from a sparse graph
$l_{ij}$ Euclidean distance between nodes $(i,j)$
$l_{k}$ Travel time of arc $k\in A$
Distance measures
$l_{\min}[i]$ Shortest-path distance from the depot to node $i$ (Dijkstra)
$l_{\max}[i]$ Longest-path estimate from the depot to node $i$ (Bellman-Ford-based)
Time windows
$a_{i}$ $\forall i\in V_{r}$: earliest service time generated as combinations of $l_{\min}[i]$ and $l_{\max}[i]$, with an additional random perturbation
@@
$b_{i}$ $\forall i\in V_{r}$: $b_{i}=a_{i}+w_{i}$, where $w_{i}$ is a random offset
$\forall i\in V\setminus V_{r}$: $b_{i}=10^{5}$
$s_{i}$ $\forall i\in V_{r}$: $s_{i}\in\{5,10\}$
@@
Demands and capacity
$d_{i}$ $\forall i\in V_{r}$: 50% pickups $[10,20]$, 50% deliveries [5,1][-5,-1]
@@
$Q$ Vehicle capacity: $Q=\sum_{i\in V}|d_{i}|+\theta$, with $\theta$ a small random offset
Auxiliary parameter
$M$ Big-M constant: $M=\max_{i}b_{i}+\max_{i}s_{i}+\max_{(i,j)\in A}l_{ij}$
To enhance the replicability of this study, all generated instances and the instance generation code are publicly available at [10].